The sum of the ages of Mr. Smith, Ms. Smith and their son, Tommy Smith, is 74. Ten years ago, this sum was 47. How old is Ms. Smith now if Mr. Smith is 28 years older than Tommy?
can you explain and show how you did the problem?
thanks

Answer by ankor@dixie-net.com(22740)

Write an equation for each statement:
:
The sum of the ages of Mr. Smith, Ms. Smith and their son, Tommy Smith, is 74.
a + b + c = 74
:
Ten years ago, this sum was 47.
(a-10) + (b-10) + (c-10) = 47
a + b + c - 30 = 47
a + b + c = 47 + 30
a + b + c = 77, but how could this be when above the sum is 74